OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, −4) and a directrix of y = 4. can someone help me with this pre test question

OpenStudy (anonymous):

a - f(x) = −16x2 b - f(x) = 16x2 c - f(x) = −one sixteenth x2 d - f(x) = one sixteenthx2

OpenStudy (perl):

i would start by drawing a picture, u can make a rough graph using the information given

OpenStudy (anonymous):

how do i do that on here

OpenStudy (anonymous):

i will give medal to whoever gives me right answer

OpenStudy (anonymous):

My answer was c btw

OpenStudy (anonymous):

its c

OpenStudy (anonymous):

thanks ! im glad i was right

OpenStudy (anonymous):

step 1: H=0, K+P = -4 step 2: y=k−p=-4 step 3: k+p = -4, k-p = 4